hans is 3 years younger than his brother emil. in 8 years the sum of their ages will be 61. how old is hans plz i need this in 30 mins plz help me im stuck
Answer by Maths68(1474)


Let
Present age of Hans = h
Present age of Emil = e

Hans is 3 years younger than his brother Emil.
h=e-3.................(1)

In 8 years
Hans will be = h+8
Emil will be = e+8

Then

The sum of their ages will be 61
(h+8)+(e+8)=61
h+8+e+8=61
h+e+16=61
h+e=61-16
h+e=45..................(2)

Put the value of h from (1) to (2)
h+e=45..................(2)
e-3+e=45
2e-3=45
2e=45+3
2e=48
2e/2=48/2
e=24

Put the value of e in (1)
h=e-3.................(1)
h=24-3
h=21

Present age of Hans = h = 21 years old
Present age of Emil = e = 24 years old